# Benchmark of segmentation techniques for pelvic fracture in CT and X-ray: Summary of the PENGWIN 2024 challenge

### v1

@@ 4.6 False-positive analysis
In our evaluation framework, false-positive (FP) fragments were not explicitly penalized at the instance level, but were instead indirectly reflected through the anatomical segmentation metrics. To better understand the prevalence and cause of FPs, we additionally computed the number of unmatched predicted fragments for each participating algorithm with respect to the ground truth, regardless of its size or shape. This analysis provides insight into over-segmentation tendencies and complements the primary evaluation metrics. The results are summarized in Table 7.
In the CT task, team SMILE achieved the lowest FP rate, attributable to its coarse-to-fine anatomical segmentation strategy, which effectively excluded non-pelvic bony structures, and the use of a primary-secondary representation scheme that tended to under-segment small fragments rather than over-segment them. The algorithm developed by team Sano exhibited a higher FP rate due to its tendency to over-segment small fragments. However, a portion of these fragments actually appeared in plausible locations—such as regions with thin, deformed cortical bone—and could be eliminated through post-processing with size-based filtering when needed. Other sources of FP predictions included erroneous anatomical segmentation that mistakenly included patients’ hands or the CT table, as well as over-detection in regions where fractures commonly occur.
In the X-ray task, the primary-secondary semantic segmentation approach adopted by team SMILE and team MedIG resulted in fewer FPs compared to the instance segmentation strategies employed by team LME and team luckyjing. While instance-based methods hold potential for more precise identification of small fragments, they tend to be less robust in practice.
4.7 Ranking robustness
@@
To evaluate the robustness of the final team rankings against variations in the test set, we performed a bootstrap analysis using 1,000 randomly resampled replacements (Wiesenfarth et al., 2021). For each bootstrap sample, the final ranks were recalculated following the same evaluation protocol. We then calculated the Kendall rank correlation coefficient $\tau$ between each bootstrap-derived ranking and the original ranking computed on the full test set, and visualized the ranking variability of each team using a rank stability plot.
@@
4.8 Ensemble of participating algorithms
We investigated the feasibility of ensembling multiple segmentation algorithms to achieve improved performance and robustness.
In the CT task, to obtain a consensus segmentation from multiple team submissions, we developed a graph-based ensemble strategy tailored for instance-level fracture segmentation. This approach was applied independently to each of the three anatomical regions and incorporates spatial overlap analysis, voting-based fusion, and standardized label assignment.
Specifically, for each test case, we first extracted all candidate fragment masks from the submissions of the top five teams. Candidate fragments were grouped by anatomical region using predefined label ranges. To identify similar fragments across different submissions, we constructed a pairwise similarity graph where each node corresponds to a fragment and edges connect pairs with an IoU greater than 0.5. Clusters were then identified as the connected components of this graph, each representing a consensus fragment candidate. Within each cluster, a majority voting scheme was applied to the set of masks to generate a single fused segmentation. Clusters with fragment size less then 500 mm3 were excluded. To address overlapping masks, fragments were written in order of size, allowing smaller fragments to overwrite larger ones and thus better preserve detail in thin regions.
As shown in the last row of Table 5, the method achieved superior fracture segmentation performance across all three metrics compared to the participating algorithms, while attaining anatomical segmentation performance comparable to that of the top-performing method.
In the X-ray task, the projection nature of the images introduces ambiguity, making it difficult to determine whether two overlapping fragments from different algorithms represent the same structure or distinct fragments that overlap in the projection. To mitigate this, we adopted a conservative merging strategy of the four teams: all candidate fragments were initially preserved, and only those with a high degree of spatial agreement—defined by an IoU greater than 0.9—were merged. While this approach helped avoid erroneous merging of distinct structures, it also increased the risk of including FPs or redundant, highly similar fragments in the final output.
As shown in the last row of Table 6, the ensemble method leveraged the diversity of participating algorithms to achieve improved fragment-wise segmentation performance. However, this gain came at the cost of a notable drop in anatomical segmentation accuracy, due to an increased number of FP predictions.
5 Discussion
5.1 The overall scheme for fragment segmentation
@@
In the CT task, two-stage approaches consisting of anatomical segmentation followed by fragment segmentation were widely adopted by the participants. This two-stage strategy was chosen for multiple reasons: (1) The sacroiliac joint, whether dislocated due to trauma or not, exhibits distinct characteristics compared to fracture surfaces within individual pelvic bones, making it difficult to effectively address both in a single-step segmentation approach; (2) The effectiveness of this scheme has been previously validated in the FracSegNet study, where bone region extraction simplified the task; (3) Pre-trained large models, such as TotalSegmentator and the CTPelvic1K baseline model, could be directly utilized to extract the bone regions; and (4) The evaluation metric design explicitly required accurate anatomical segmentation as a distinct criterion, ensuring alignment with the requirements of subsequent surgical planning procedures (Liu et al., 2025). We computed the correlation between IoU-A and IoU-F in the CT task across all participating teams. A strong linear relationship was observed, with Pearson’s correlation coefficient of 0.870 (p = 0.0005), suggesting that accurate anatomical segmentation serves as a critical foundation for successful fracture segmentation.
The primary distinction among the top teams lay in their fracture segmentation strategies in the second stage. To convert the instance segmentation problem, which involves a variable number of fragments, into a more tractable semantic segmentation problem, some teams adopted a primary-secondary representation, while others employed a boundary-core representation to isolate individual fragment instances.
The primary-secondary representation can be challenging when the definition of the ”primary” fragment is ambiguous or when multiple secondary fragments collide in adjacent regions, a scenario frequently observed in anterior pelvic ring fractures. In contrast, the boundary-core representation is well-suited for handling such cases but depends heavily on precise boundary segmentation. Small errors in boundary prediction can lead to significant connectivity changes, which is a potential concern when dealing with incomplete fractures.
The solutions in the X-ray task exhibited greater methodological diversity compared to the CT task. While some teams employed two-stage approaches similar to those used in the CT task, others opted for more direct methods, leveraging advances in 2D image detection and segmentation techniques, such as Mask R-CNN. These approaches, which have demonstrated strong performance across various applications, were adapted to accommodate multi-class overlapping labels, addressing the unique challenges of the X-ray segmentation task.
All of the four methods summarized in this task focused on fragment labels rather than fracture lines, as overlapping fragments remain inseparable due to the projection nature of X-ray imaging, even when the fracture regions are removed. The primary-secondary representation, despite being adopted by both the first- and second-place teams, exhibited a notable limitation: small fragments classified as secondary were more frequently connected compared to the CT images, making them difficult to separate through post-processing. This challenge posed a significant drawback in accurately delineating individual small fragments with semantic segmentation approaches.
The results of X-ray segmentation algorithms proved insufficient to reliably support clinical decision-making or subsequent tasks. Several factors contributed to this limitation: (1) A single bone fragment may appear as multiple disconnected parts in the X-ray due to the limited field of view; and (2) The overlap of fragments with other tissues significantly compromises the delineation of fracture lines and hinders fragment separation.
To overcome these challenges, incorporating prior knowledge of the fragments is essential. One potential approach is to utilize multiple X-ray images from different views or targeted regions, as overlapping fragments in one projection may appear as distinct regions in a near-orthogonal plane, thereby improving segmentation accuracy with the mutually complementary information. Another promising direction involves integrating preoperative CT segmentation into the process. Given the availability of a patient’s CT scan, 2D/3D registration can be employed to estimate the approximate positions of fragments (Grupp et al., 2019). However, additional efforts are required to account for the exact C-arm geometry, which is typically unknown beforehand, and to address fragment position changes that occur between CT and X-ray acquisition, both before and during surgery. Although these methodologies extend beyond the current scope of the PENGWIN challenge, they hold significant potential for translating fracture segmentation techniques into clinical practice.
5.2 Limitations and future directions of PENGWIN
5.2.1 Dataset
@@
Compared to related challenges organized at major medical imaging conferences, where the number of CT samples typically ranges from 160 to 1,400, the PENGWIN dataset is relatively small in size. This limitation is primarily due to the rarity of pelvic fracture CT data, as pelvic fractures account for only 3% of all fractures (Hu et al., 2023).
@@
Additionally, while the X-ray dataset closely resembles real clinical data due to the DeepDRR simulation approach, it remains synthetic rather than derived from real patient imaging. Although DeepDRR facilitates high-fidelity X-ray generation, enabling this prospective study with a large annotated dataset, the absence of real C-arm X-ray images raises concerns regarding its direct clinical translation.
For future iterations of the PENGWIN challenge, we plan to substantially expand the fractured CT dataset, increasing the number of cases to 300–500 scans. Moreover, we aim to incorporate real clinical C-arm X-ray images in test set and annotate them using 2D/3D registration approaches, aligning them with corresponding CT scans to ensure ground-truth accuracy and facilitate clinically relevant segmentation tasks.
5.2.2 Uncertainty in annotation
@@
As discussed in Sec. 4.4, the annotation process involved inherent uncertainty, requiring expert annotators to determine whether an incomplete fracture (partially fractured but non-displaced) should be classified as two separate fragments or a single piece. This decision was primarily based on their understanding of the subsequent surgical process, particularly whether the two segments could be manipulated as a single unit during reduction without undergoing significant deformation.
@@
5.2.3 Challenge logistics
Finally, we summarize several logistical limitations identified during the challenge and the post-challenge participant survey, which should be addressed in future iterations and serve as lessons for other challenge organizers to consider.
First, although the dataset and preliminary validation phase were released well in advance, the final test phase deadline was extended by three days due to multiple participant requests. To maintain fairness, teams that had already submitted their algorithms were given an opportunity to update their submissions. However, such extensions should be avoided in future iterations to uphold the rigor and integrity of the challenge timeline.
@@ 5.2.3 Challenge logistics
Lastly, we encountered challenges in retrospectively collecting algorithm descriptions from participants after the challenge had ended, particularly from lower-performing teams. To address this, future iterations will require participants to submit a technical description of their method along with their final submission, ensuring that essential technical details are documented in a timely manner.
6 Conclusion
In this paper, we have presented the PENGWIN challenge on pelvic fracture segmentation in CT and X-ray images, organized as part of MICCAI 2024. We provided a comprehensive review of related challenges and studies, a detailed overview of the challenge setup, dataset, participant contributions, and performance analysis.
In the CT task, the winning team achieved a high level of accuracy, with an average fragment-wise IoU of 0.930. In the X-ray task, the best-performing algorithm attained an IoU of 0.774, which, while promising, remains insufficient for reliably supporting clinical applications.
Beyond quantitative performance, the participating algorithms demonstrated great methodological diversity, offering valuable insights for future research in this emerging field. The various approaches and strategies employed have been thoroughly discussed in this paper. While the focus of this challenge was on pelvic fractures, we believe that many of the developed methodologies and key findings will generalize to other types of fractures. The pelvis presents one of the most complex bone anatomies, making it an ideal test case for evaluating and refining fracture segmentation techniques.
For future iterations of the PENGWIN challenge, we aim to expand the challenge dataset and explore interactive segmentation approaches that integrate human decision-making with task-relevant information, thereby enhancing model performance in handling uncertainty and improving clinical applicability.

### v2

@@ IV-F False-Positive Analysis
In our evaluation framework, false-positive (FP) fragments were not explicitly penalized at the instance level, but were instead indirectly reflected through the anatomical segmentation metrics. To better understand the prevalence and cause of FPs, we additionally computed the number of unmatched predicted fragments for each participating algorithm with respect to the ground truth, regardless of its size or shape. This analysis provides insight into over-segmentation tendencies and complements the primary evaluation metrics. The results are summarized in Table VI.
In the CT task, SMILE achieved the lowest FP rate, attributable to its coarse-to-fine anatomical segmentation strategy, which effectively excluded non-pelvic bony structures, and the use of a primary-secondary representation scheme that tended to under-segment small fragments rather than over-segment them. The algorithm developed by Sano exhibited a higher FP rate due to its tendency to over-segment small fragments. However, a portion of these fragments actually appeared in plausible locations—such as regions with thin, deformed cortical bone—and could be eliminated through post-processing with size-based filtering. Other sources of FP predictions included erroneous anatomical segmentation that mistakenly included patients’ hands or the CT table, as well as over-detection in regions where fractures commonly occur.
In the X-ray task, the primary-secondary semantic segmentation approach adopted by SMILE and MedIG resulted in fewer FPs compared to the instance segmentation strategies employed by LME and luckyjing. While instance-based methods hold potential for more precise identification of small fragments, they tend to be less robust in practice.
IV-G Ranking Robustness
@@
To evaluate the robustness of the final team rankings against variations in the test set, we performed a bootstrap analysis using 1,000 randomly resampled replacements [wiesenfarth2021methods]. For each bootstrap sample, the final ranks were recalculated following the same evaluation protocol. We then calculated the Kendall rank correlation coefficient $\tau$ between each bootstrap-derived ranking and the original ranking computed on the full test set, and visualized the ranking variability of each team using a rank stability plot.
@@
IV-H Example Failure Cases
Figure 11: Example failure modes. The red and yellow boxes indicate false-negative and false-positive regions, respectively.
Fig. 11 presents representative failure cases from both the CT and X-ray tasks, illustrating common error patterns observed across participating algorithms.
In CT images, false negatives (missed fragments) predominantly occur in cases involving very small fragments, particularly near the sacroiliac joint, where anatomical segmentation may under-segment the surrounding bone and obscure fragment boundaries (see Sec. IV-F). Conversely, false positives (spurious fragments) often arise in incomplete or non-displaced fractures, where subtle cortical irregularities may be over-interpreted as separable fragments (see Sec. V-C2).
In X-ray images, both false negatives and false positives are primarily driven by occlusion and projection overlap. For example, superimposed anterior-ring fragments in inlet projections can render multiple fragments indistinguishable, leading to missed detections. Similarly, high-frequency edges introduced by bowel gas or soft-tissue–bone overlap may be mistakenly identified as fracture boundaries. These examples highlight the intrinsic ambiguity of 2D projections and the difficulty of recovering 3D fragment topology from overlapping structures (see Sec. V-A2).
V Discussion
V-A The Overall Scheme for Fragment Segmentation
V-A1 CT Segmentation
@@
In the CT task, two-stage approaches consisting of anatomical segmentation followed by fragment segmentation were widely adopted by the participants. This two-stage strategy was chosen for multiple reasons: (1) The sacroiliac joint, whether dislocated due to trauma or not, exhibits distinct characteristics compared to fracture surfaces within individual pelvic bones, making it difficult to effectively address both in a single-step segmentation approach; (2) The effectiveness of this scheme has been previously validated in the FracSegNet study, where bone region extraction simplified the task; (3) Pre-trained large models, such as TotalSegmentator and the CTPelvic1K baseline model, could be directly utilized to extract the bone regions; and (4) The evaluation metric design explicitly required accurate anatomical segmentation as a distinct criterion, ensuring alignment with the requirements of subsequent surgical planning procedures [liu2025preoperative]. We computed the correlation between IoU-A and IoU-F in the CT task across all participating teams. A strong linear relationship was observed, with Pearson’s correlation coefficient of 0.870 (p = 0.0005), suggesting that accurate anatomical segmentation serves as a critical foundation for fracture segmentation.
A key methodological divergence among the top-performing teams in the CT task was their approach to instance-level fracture segmentation in the second stage of the pipeline. Given the challenge of predicting a variable number of fragments, teams reformulated the problem into a semantic segmentation task using one of two main strategies: primary-secondary labeling or boundary-core decomposition. The primary-secondary strategy assigns a single dominant (primary) fragment label to each bone and groups the remaining pieces as secondary fragments, whereas the boundary-core approach segments the core regions of fragments and delineates their boundaries to enable instance separation.
Each strategy offers trade-offs. The primary-secondary method is conceptually simple and well-suited for most fracture patterns, but its effectiveness diminishes in cases where the designation of a primary fragment is ambiguous—such as bilateral fractures or when secondary fragments are of comparable size. Additionally, secondary fragments that are spatially adjacent may become fused during post-processing, reducing instance accuracy. Conversely, the boundary-core approach facilitates more precise separation of contacting or overlapping fragments, which is especially beneficial in complex fracture regions like the anterior pelvic ring. However, this method is sensitive to small boundary segmentation errors, which may disconnect or erroneously merge fragments, particularly in the presence of hairline or incomplete fractures.
The ABBC formulation adopted by the winning MIC-DKFZ team demonstrated a clear performance advantage, particularly in complex multi-fragment cases. By explicitly modeling fragment separability, the method first predicts the core region of each fragment and then learns a thin boundary representation that delineates where fragments separate. This reduces dependence on heuristic post-processing steps and substantially lowers the risk of fragment merging that commonly arises in primary–secondary pipelines. While the boundary–core approach introduces increased sensitivity to local boundary mis-segmentation and requires careful control of splitting behavior, its strong ability to robustly separate fragments that are in close apposition appears to be a key factor behind the superior IoU and surface distance metrics observed. A broader takeaway is that explicitly modeling fracture boundaries—rather than relying solely on semantic labeling—offers a principled advantage in anatomies where structural continuity is disrupted in complex and clinically variable ways. Moreover, because the formulation does not depend on bone shape priors, it can be directly applied to other fracture anatomies without requiring changes to the labeling policy.
V-A2 X-Ray Segmentation
In contrast to the relatively consistent two-stage pipelines in the CT task, the X-ray task showcased greater methodological diversity, particularly in the overall segmentation schemes adopted by participating teams. These differences reflect the inherent challenges of 2D projection imaging, where occlusion and overlapping structures complicate the delineation of fracture fragments.
As shown in Fig. 5 and Table III, the most notable distinction among top-performing teams lies in how they structured their overall approach. Some teams, such as SMILE and LME, employed a two-stage strategy—first isolating the pelvic bones, followed by fragment segmentation. This separation can simplify the learning problem for the fragment segmentation module by removing irrelevant background and anatomical clutter. However, it may introduce compounding errors: inaccuracies in the anatomical segmentation stage could propagate and affect downstream fragment predictions.
Other teams, like MedIG and luckyjing, opted for a single-stage strategy that directly outputs either semantic or instance-level labels from the original input image. Notably, MedIG used three independent networks for segmenting the three bone regions separately, whereas others used a single network with multi-label prediction. Using separate networks allows specialization and potentially better accuracy for each bone, especially when the visual characteristics or fragment patterns differ, but at the cost of computational efficiency. In contrast, using a single network with shared backbone and multi-class output offers better efficiency and enforces global anatomical coherence, but may suffer from performance degradation in locations that are underrepresented or visually ambiguous.
Efficiency considerations are especially relevant for the X-ray task, where segmentation may be required in time-sensitive clinical settings such as emergency trauma care or intra-operative fluoroscopic guidance. While we report container runtimes as an efficiency proxy, these measurements include substantial overhead from environment initialization and I/O, and therefore do not reflect pure model inference time. Additionally, several teams employed model ensembling and repeated inference strategies to maximize accuracy within the runtime limit, trading computational efficiency for performance. Notably, the Mask R-CNN-based pipelines demonstrated faster inference on 2D images and, with appropriate pruning or lightweight backbone substitution, may be adaptable for near real-time use. In contrast, two-stage anatomical-then-fragment pipelines, though accurate, incur higher computational load and are less compatible with fast-turnaround workflows.
Across all methods, the target remained consistent: fragment-level labeling, rather than fracture-line segmentation, as the latter is confounded by projection overlap in X-rays. Teams that employed the primary-secondary labeling, such as SMILE, faced challenges in separating small fragments, which tended to appear fused in the projection space. While instance segmentation models like Mask R-CNN (LME and luckyjing) offer potential for explicit fragment separation, they remain limited by the lack of depth cues and can struggle with over-segmentation or missing small fragments.
These varying design choices underline the difficulty of the X-ray segmentation task and point to important trade-offs between modularity, anatomical specificity, and model generalization. Despite the similar task formulation, the segmentation performance on X-ray images was notably lower than on CT. For instance, the top-performing method in the CT task achieved a IoU-F of 0.93, whereas the top method in the X-ray task only reached 0.77. From a clinical perspective, such a discrepancy has substantial implications: in surgical planning or intraoperative guidance, insufficiently separated or misidentified fragments may lead to incorrect assessments of fracture morphology or fragment displacement, ultimately affecting surgical strategy.
The reduced accuracy in X-ray segmentation arises from intrinsic limitations of projection imaging: (1) A single 3D fragment may appear disconnected or distorted in 2D projections due to occlusion or truncation by the limited field of view; and (2) Overlapping anatomical structures frequently obscure fracture lines, especially in complex or comminuted injuries, making instance-level identification extremely challenging. These factors collectively compromise the ability of current semantic and instance segmentation methods to distinguish fragments robustly in X-rays.
To bridge this performance and applicability gap, future efforts could incorporate prior anatomical knowledge or leverage additional imaging information. One strategy is the acquisition of multiple views or targeted X-ray angles; complementary perspectives could help disambiguate fragment boundaries that are occluded in a single view. Another direction is to integrate preoperative CT segmentation via 2D/3D registration [grupp2019pose], providing a rough anatomical prior for localizing fragments in intraoperative X-rays. However, practical deployment of this approach faces challenges, including estimation of C-arm pose and compensating for changes in patient positioning or fragment displacement between scans.
Although these strategies exceed the scope of the current challenge, they reflect the necessary steps toward translating X-ray segmentation into reliable clinical tools—particularly in image-guided orthopedic procedures, where real-time, accurate fragment localization is critical.
V-B Impact of Segmentation on Reduction Planning
Figure 12: Example reduction planning result using different segmentation approaches. The anatomical alignment is affected by the separability of fragments.
To demonstrate the clinical implications of segmentation quality, particularly its impact on downstream surgical planning, we conducted a comparative analysis using the FracFormer reduction planning framework [yibulayimu2025fracformer]. We evaluated reduction outcomes derived from two representative CT segmentation strategies: the boundary–core decomposition used by MIC-DKFZ and the primary–secondary labeling scheme used by SMILE.
As illustrated in Fig. 12, differences in fragment separability directly propagate into the reduction stage. When small fragments are not correctly isolated—an issue more commonly observed in primary–secondary pipelines—fused or misidentified fragments introduce incorrect geometric constraints during alignment, causing the estimated pose to drift away from the desired anatomical configuration. This observation is consistent with prior findings showing that manually separating small fragments on automated segmentation results can improve reduction pose estimation [liu2025preoperative]. In the example shown, the boundary–core segmentation produced a chamfer distance of 1.695 mm, compared with 2.097 mm for the primary–secondary result. Although both errors fall within clinically acceptable ranges, this experiment underscores a key insight: instance-level separability—rather than volumetric overlap alone—is critical for downstream tasks that depend on accurate geometric reconstruction, such as reduction pose/trajectory estimation and fixation strategy design.
V-C Limitations and Future Directions
V-C1 Dataset
@@
Compared to related challenges organized at major medical imaging conferences, where the number of CT samples typically ranges from 160 to 1,400, the PENGWIN dataset is relatively small in size. This limitation is primarily due to the rarity of pelvic fracture CT data, as pelvic fractures account for only 3% of all fractures [hu2023epidemiology]. Additionally, all cases were collected from clinical centers in China, representing a relatively homogeneous patient population. This may limit the generalizability of trained models across different demographic groups and imaging practices.
@@
While the X-ray dataset closely resembles real clinical data due to the DeepDRR simulation approach, it remains synthetic rather than derived from real patient imaging. Although DeepDRR facilitates high-fidelity X-ray generation, enabling this prospective study with a large annotated dataset, the absence of real C-arm X-ray images raises concerns regarding its direct clinical translation. Due to the 2D projection nature, inter-observer variability was not evaluated in the X-ray task, as consistent delineation of overlapping fragments was not achievable even among trained annotators. In addition, multiple X-ray projections were generated from each CT scan, and individual projections were treated as independent test samples during evaluation. While variation in projection angle leads to meaningful differences in occlusion and visibility patterns, multiple projections per subject introduce repeated measures and reduce strict statistical independence in the test set.
For future iterations of PENGWIN, we plan to substantially expand the fractured CT dataset, increasing the number of cases to 300–500 scans. Furthermore, we aim to incorporate real intraoperative C-arm X-ray images into the test set. Since fragment-level annotation in X-ray images is inherently challenging, these images will be labeled using a 2D/3D registration pipeline to align them with preoperative CT segmentations, ensuring anatomical consistency and clinically meaningful ground truth. This will also allow us to evaluate sim-to-real generalization performance, building on recent findings showing that DeepDRR-generated synthetic data can support transfer to real clinical X-rays [gao2023synthetic].
V-C2 Inaccuracy in Annotation
The initial anatomical segmentation used to support the annotation workflow was generated by a trained automatic model and subsequently refined by expert annotators. While this approach significantly reduced manual workload and improved structural consistency, it may introduce subtle bias from the model’s prior assumptions.
@@ V-C2 Inaccuracy in Annotation
As discussed in Sec. IV-D, the refinement process involved inherent uncertainty, requiring annotators to determine whether an incomplete fracture should be classified as two separate fragments or a single piece. This decision was primarily based on their understanding of the subsequent surgical process, particularly whether the two segments could be manipulated as a single unit during reduction without undergoing significant deformation.
@@
V-D Recommendations for Future Challenges
We summarize several practical lessons identified during the challenge and the post-challenge participant survey, which should be addressed in future iterations and may serve as useful guidance for other challenge organizers.
First, although the dataset and preliminary validation phase were released well in advance, the final test phase deadline was extended by three days due to multiple participant requests. To maintain fairness, teams that had already submitted were given an opportunity to update their submissions. However, such extensions should be avoided in future iterations to uphold the rigor and integrity of the challenge timeline.
@@ V-D Recommendations for Future Challenges
Lastly, we encountered challenges in retrospectively collecting algorithm descriptions from participants after the challenge had ended, particularly from lower-performing teams. To address this, future challenges may require participants to submit a technical description of their method along with their final submission, ensuring that essential technical details are documented in a timely manner.
VI Conclusion
We have presented the PENGWIN challenge on pelvic fracture segmentation in CT and X-ray images, organized as part of MICCAI 2024. We provided a comprehensive review of related challenges and studies, a detailed overview of the challenge setup, dataset, participant contributions, and performance analysis.
In the CT task, the winning team achieved a high level of accuracy, with an average fragment-wise IoU of 0.930. In the X-ray task, although the top-performing method reached an IoU of 0.774, this level of accuracy remains insufficient for reliable clinical use. Future work will incorporate evaluation on real C-arm X-ray datasets and explore multi-view or CT-informed approaches to mitigate projection ambiguity.
Beyond quantitative performance, the participating algorithms demonstrated great methodological diversity, offering valuable insights for future research in this emerging field. The various approaches and strategies employed have been thoroughly discussed in this paper. While the focus of this challenge was on pelvic fractures, we believe that many of the developed methodologies and key findings will generalize to other types of fractures.